find the indicated outputs for f(x)=3x(2)-4x
f(0)=
f(-1)=
f(2)=

Answer by checkley77(12844)

f(x)=3x(2)-4x
f(x)=6x-4x
f(x)=2x
f(0)=2*0=0 ans.
f(0)=0
f(-1)=2-1=1 ans.
f(-1)=1
f(2)=2*2=4 ans.
f(2)=4